OpenStudy (anonymous):

1+log3(5+x)=2log3(1−x ) not just (log3) the 3 should be below, bit down the log

OpenStudy (anonymous):

a bit down

OpenStudy (alfie):

\[1+\log_3(5+x) = 2\log_3(1-x) \rightarrow \log_33 +\log_3(5+x) = 2\log_3(1-x)\] \[\log_33(5+x) = \log_3(1-x)^2\] 15+3x = (1-x)^2 Solve you get two solutions... x = -2 x = 7. The only one acceptable is x = -2 because of the domain : 5+x > 0 => x > - 5 -x > -1 => x < 1 -5<x<1 So x = -2